major axis 20 units long and parallel to y-axis minor axis 6 units long center at (4,2)
Answer by lwsshak3(11628)

major axis 20 units long and parallel to y-axis minor axis 6 units long center at (4,2)
**
Standard form of equation for ellipse with vertical major axis: (x-h)^2/b^2+(y-k)^2/a^2, a>b, (h,k) being the (x,y) coordinates of the center.
For given ellipse:
center: (4,2)
length of vertical major axis=20=2a
a=10
a^2=100
..
length of minor axis=6=2b
b=3
b^2=9
..
Equation of given ellipse: (x-4)^2/9+(y-2)^2/100=1